ROTATIONAL ANALYSIS OF THE A^{2}Π − X^{2}Σ^{+} AND B^{2}Σ^{+} − X^{2}Σ^{+} TRANSITIONS OF YbBr USING LASER EXCITATION AND RESOLVED FLUORESCENCE

\underline{CAMERON S. DICKINSON} and JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, NS, Canada B3H 4J3; COLAN LINTON, Department of Physics, University of New Brunswick, P.O. Box 4400, Fredericton, New Brunswick, Canada E3B 5A3.

High-resolution laser excitation spectra have been recorded for the first time for YbBr. Data for the 0-0 and 1-0 bands of the A^{2}Π − X^{2}Σ^{+} system and the 0-0, 1-0 and 1-1 bands of the B^{2}Σ^{+} − X^{2}Σ^{+} system were obtained using selective detection of fluorescence. The lowest vibrational levels of the A^{2}Π_{3/2} and B^{2}Σ^{+} states are shown to be separated by ≈ 400 cm^{−1}, and exhibit the characteristics of a mutually perturbing pair. Through observation of symmetry forbidden transitions in the B^{2}Σ^{+} − X^{2}Σ^{+} system, mixing of the A and B states is apparent. A resolved fluorescence study has also been undertaken by observing the 0-1, 0-2 and 0-3 bands in the A^{2}Π_{1/2} − X^{2}Σ^{+} system. A least-squares merged fit of the estimated parameters from individual bands was performed to remove redundancies, the results of which will be presented.